Minute of KAGRA commissioning meeting
- Date: 2018/7/24 11AM~12PM
- Attendance: Miyakawa, Izumi, Ohishi
- eZuce: Yamamoto, Nakano, Shoda, Suzuki
- EOM and FI are being tested.
- EOM: TF and modulation index were measured.
- Fiber isolate reached and is being tested.
- Installation to KAGRA started.
- MZ alignment is being proceeded, and MC has been already locked with single arm of MZ.
- Mode matching ratio is still bad but it will be fixed soon.
- ISS loop locked again. OLG or RIM was measured, and they are being analyzed.
- On 7/9, we started take IMMT out from the chamber.
- All the replacement of magnetized screws is done by the end of last week.
- Mirror position (distance between mirror and coile) was shifted a ~mm, probably caused by tilt of optical.
- TF will be measured soon.
- Old circuits are still used in TAMA era, and all the noise floor are high.
To fix BS GeoPhone, we should open the chamber.
TMS GeoPhone looks fine, but some strange coupling can be seen in the TF measurement.
- Feedback bandwidth is limited ~10Hz.
- DAC is almost out of stock. We should think to start using old DAC used in CLIO.
- Next meeting: 7/31.